# Analysis of Discrete Modern Hopfield Networks in Open Quantum System

### v1

@@ IV Conclusion
In this study, we propose an open quantum discrete modern Hopfield, a generalization of the open quantum Hopfield network, and determine its time evolution and the phase diagram. Unlike the conventional Hopfield model ($x=2$), the origin ($M_{Z}=0$) is always stable in our the model ($x=4$), and the region with an LC is smaller. Each phase is determined by combining the analytical calculation of the number of fixed points with numerical analysis of LC’s existence. Although our model qualitatively differs from the open quantum Hopfield model ($x=2$), the behavior of $x>2$ not changes significantly.
As the classical modern Hopfield networks have broad applications and large capacity, we expect that our quantum variant will offer similar advantages. However, the strong nonlinearity makes the interactions multivariate, making the analytical calculation of the capacity challenging. The memory capacity for the open quantum Hopfield model is calculated through the dilution of the network and the cumulant expansion; howevwr, the applicability of these tools to our model remains unclear.
Key subclasses of classical modern Hopfield networks include transformer and diffusion models, defined as continuous variable models. To consider the quantum analog of the building blocks of modern neural networks, future studies will extend our model to continuous variable systems. This generalization will be addressed as an important future problem in subsequent studies.
@@ Acknowledgements.
T. K. acknowledges support from JST SPRING, Grant Number JPMJSP2125 and would like to take this opportunity to thank the “THERS Make New Standards Program for the Next Generation Researchers”. K. K. acknowledges support from JSPS Grant-in-Aid for Early-Career Scientists, No. 22K13972; from MEXT-JSPS Grant-in-Aid for Transformative Research Areas (B), No. 24H00829.

### v2

@@ IV Conclusion
In this study, we propose an open quantum discrete modern Hopfield network, a generalization of the open quantum Hopfield network, and determine its time evolution and phase diagram. Unlike the conventional Hopfield model ($x=2$), the origin ($M_{Z}=0$) is always stable in our model ($x=4$), and the region with the LC is smaller. Each phase is determined by combining the analytical calculation of the number of fixed points with numerical analysis of the LC’s existence. Although our model qualitatively differs from the open quantum Hopfield model ($x=2$), the behavior of $x>2$ does not change significantly.
As the classical modern Hopfield networks have broad applications and large capacity, we expect that our quantum variant will offer similar advantages. However, the strong nonlinearity makes the interactions multivariate, making the analytical calculation of the capacity challenging. The memory capacity for the open quantum Hopfield model is calculated through the dilution of the network and cumulant expansion; however, the applicability of these tools to our model remains unclear.
As future directions, we propose two promising extensions of our model that aim to enhance its realism and applicability. First, one can introduce interactions between each node and its local environment to model the effects of noise. This is particularly relevant for realistic implementations, where local environmental interactions are unavoidable. Studying how such noise influences the system’s behavior could offer valuable insights into the model’s robustness under practical conditions. Second, a natural extension is to explore time-dependent coupling to the system, particularly in the context of quantum annealing. Since our framework ultimately aims at energy minimization, implementing an annealing schedule, where the temperature is gradually lowered over time, would be a compelling direction. In this context, identifying effective annealing protocols becomes a central issue, especially for practical use cases.
In addition to these discrete-variable models, another important direction is to generalize the framework to continuous variable systems. Key subclasses of classical modern Hopfield networks, such as transformer and diffusion models, are formulated in this way. Investigating quantum analogs of these continuous architectures remains a challenging and significant goal for future research.
@@ Acknowledgements.
T.K. acknowledges support from MEXT Q-LEAP JPMXS0120319794 and MEXT Q-LEAP JPMXS0120351339; from JST SPRING, Grant No. JPMJSP2125, and would like to thank the “THERS Make New Standards Program for the Next Generation Researchers.” K.K. acknowledges support from JSPS KAKENHI Grant-in-Aid for Early-Career Scientists (Grant No. JP22K13972), Grant-in-Aid for Transformative Research Areas (B) (Grant No. JP24H00829), and Grant-in-Aid for Scientific Research (C) (Grant No. JP23K17668).